# Towards a Science of Scaling Agent Systems

### v1

@@ 4.3 Scaling principles
Mixed-Effects Model Achieves 51.3% Cross-Validated Variance Explanation.
We fit a scaling principle to all 180 configurations that relates agentic system performance to four categories of predictors: 1) base model capability (intelligence index $I$), 2) system configuration (agent count $n_{a}$), 3) task properties (tool count $T$, single-agent baseline $P_{\text{SA}}$). These are instance-level predictors capturing within-benchmark variation, distinct from the benchmark-level domain complexity $D$ defined in Appendix B, and 4) empirically measured coordination metrics from Table 5 (efficiency $E_{c}$, overhead $O\%$, error amplification $A_{e}$, message density $c$, redundancy $R$). Rather than including all possible terms, we construct the model based on specific mechanistic hypotheses.
@@
$P=\beta_{0}$ $+\beta_{1}I+\beta_{2}I^{2}+\beta_{3}\log(1+T)+\beta_{4}\log(1+n_{a})$
$+\beta_{5}\log(1+O\%)+\beta_{6}c+\beta_{7}R+\beta_{8}E_{c}+\beta_{9}\log(1+A_{e})$
$+\beta_{10}P_{\text{SA}}+\beta_{11}(I\times E_{c})+\beta_{12}(A_{e}\times P_{\text{SA}})$
$+\beta_{13}(O\%\times T)+\beta_{14}(R\times n_{a})+\beta_{15}(c\times I)$
$+\beta_{16}(E_{c}\times T)+\beta_{17}(P_{\text{SA}}\times\log(1+n_{a}))$
@@
where all predictors are standardized ($\mu=0$, $\sigma=1$) for interpretability. Log transformations are applied to right-skewed variables spanning multiple orders of magnitude ($O\%$: 0–515%; $T$: 4–16; $n_{a}$: 1–4; $A_{e}$: 1.0–17.2) to satisfy linearity assumptions. The $A_{e}\times T$ interaction retains $A_{e}$ without additional log transformation because $\log(1+A_{e})$ already appears as a main effect; including $\log(1+A_{e})\times T$ would introduce near-collinearity (VIF >8>8). Sensitivity analysis confirms qualitatively identical results under alternative specifications ($\Delta R^{2}_{\text{CV}}<0.01$). We validate model complexity through five-fold cross-validation with experiment-level holdout, which yields $R^{2}_{\text{CV}}=0.513$ ($\pm 0.052$ SD), mean absolute error MAE =0.089=0.089 ($\pm 0.011$), and root mean squared error RMSE =0.112=0.112 ($\pm 0.014$). The modest gap between training and cross-validated $R^{2}$ ($\Delta R^{2}=0.076$), combined with stable coefficient estimates across folds (coefficient of variation $18\%$ for all $|\hat{\beta}|>0.05$), indicates that the 20 parameters are justified by predictive power rather than overfitting. This model substantially outperforms simpler alternatives using only architectural labels ($R^{2}_{\text{CV}}=0.43$) or intelligence alone ($R^{2}_{\text{CV}}=0.28$), as shown in Table 3. Critically, this equation contains no dataset-specific parameters, enabling prediction on unseen task domains. Bootstrap resampling (n=1,000n=1{,}000 iterations) confirms coefficient stability (mean bootstrap SE <0.015<0.015 for all $|\hat{\beta}|>0.1$), and residual diagnostics satisfy normality (Shapiro–Wilk $p=0.412$) and homoscedasticity (Breusch–Pagan $p=0.298$), with residual standard error $\hat{\sigma}=0.118$. We evaluated regularized alternatives: Lasso (10-fold CV for $\lambda$ selection) retained 16 of 20 predictors with $R^{2}_{\text{CV}}=0.506$; Ridge achieved $R^{2}_{\text{CV}}=0.509$. Given minimal improvement and the interpretability benefits of the full model, we retain the unregularized specification.
The Efficiency-Tools Interaction Dominates Multi-Agent Performance ($\hat{\beta}=-0.330$, $p<0.001$).
@@
$\Delta P_{\text{efficiency}}=-0.330\times E_{c}\times T=\begin{cases}-2.46&\text{(single-agent, $E_{c}=0.466$)}\\ -0.39&\text{(multi-agent, $E_{c}=0.074$)}\end{cases}$
Thus, single-agent systems incur minimal efficiency penalty despite lower absolute efficiency, because the interaction magnifies the cost for architectures with many tools. Conversely, simple tasks ($T\leq 4$) show negligible efficiency effects ($|\Delta P|<0.05$), explaining why multi-agent coordination can succeed on decomposable problems. This finding contradicts the naïve hypothesis that “more agents always help with complexity”: tool-rich environments amplify the coordination tax, making simpler architectures paradoxically more effective. The effect size ($\hat{\beta}=-0.330$) is 57% larger than the next strongest interaction, establishing efficiency management as the primary bottleneck in agentic scaling.
Error Amplification Exhibits Architecture-Dependent Catastrophic Failure Modes.
Table 5 reveals dramatic variance in error amplification factors: single-agent ($A_{e}=1.0$), centralized ($A_{e}=4.4$), decentralized ($A_{e}=7.8$), hybrid ($A_{e}=5.1$), and strikingly, independent multi-agent ($A_{e}=17.2$). The scaling law captures this via the $A_{e}\times T$ interaction ($\hat{\beta}=-0.097$, $p=0.007$): errors propagate more severely in tool-rich environments. For independent architecture with $T=16$ tools, the error penalty is $0.097\times 17.2\times 16=-26.7$ (in standardized units), catastrophically degrading performance. This explains independent’s universal underperformance (mean success rate 0.370 vs. 0.466 for single-agent). The mechanism is absence of inter-agent communication: each agent operates in isolation, duplicating errors without correction opportunities. Centralized architecture, by contrast, achieves $A_{e}=4.4$ through a coordinator bottleneck that validates outputs, trading overhead (285%) for error resilience. Quantitatively, the error-overhead trade-off follows:
$\frac{\partial P}{\partial A_{e}}\bigg|_{\text{fixed }T}=\hat{\beta}_{9}\frac{1}{1+A_{e}}+\hat{\beta}_{19}T\approx-0.014-0.097T,$
indicating that error amplification becomes the dominant failure mode as $T$ increases. This formalizes the intuition that “more moving parts increase fragility,” now quantified: each additional tool amplifies error sensitivity by $\Delta A_{e}/\Delta T\approx 0.097$ (in standardized effect units).
@@
$\Delta P_{\text{overhead}}=\hat{\beta}_{5}\log(1+515)+\hat{\beta}_{13}\times 515\times 16=-1{,}159\text{ (standardized)},$
a prohibitive penalty explaining hybrid’s collapse on tool-heavy benchmarks (success rate 0.452 overall, 0.21 on workbench). The functional form implies a critical threshold:
$O\%_{\text{max}}(T)=\frac{\hat{\beta}_{5}}{\hat{\beta}_{13}T}\log(1+O\%)\approx\frac{0.032}{0.141T}\log(1+O\%),$
beyond which overhead cost exceeds any coordination benefit. For $T=16$, this threshold is $O\%\approx 150\%$, ruling out all multi-agent architectures except possibly decentralized (263%, but compensated by parallelization). Empirically, workbench confirms this prediction: decentralized (mean 0.664) outperforms centralized (0.621) despite higher overhead, due to its superior parallel efficiency. This overhead-complexity interaction constitutes the second-strongest effect ($|\beta|=0.141$), reinforcing that coordination costs are not fixed but scale super-linearly with environmental complexity.
Intelligence Exhibits Accelerating Returns via Quadratic Scaling ($\hat{\beta}_{I^{2}}=0.256$, $p=0.010$).
The inclusion of $I^{2}$ in Eq. 1 reveals non-linear capability effects: doubling intelligence yields more than double the performance gain. Specifically, for models spanning $I\in[34,66]$ (Gemini-2.0-Flash to GPT-5), the marginal benefit is:
$\frac{\partial P}{\partial I}=\hat{\beta}_{1}+2\hat{\beta}_{2}I=-0.180+0.512I$
This positive quadratic coefficient indicates accelerating returns: higher-capability models benefit disproportionately from each additional unit of intelligence, with top-quartile models (I > 60) achieving 23% higher performance than predicted by linear scaling alone. This quadratic term improves cross-validated $R^{2}$ by +0.031+0.031 ($p=0.010$), a statistically significant and operationally meaningful enhancement.
Redundancy Provides Marginal Benefit at Scale ($\hat{\beta}_{R\times n_{a}}=0.041$, $p=0.040$).
Work redundancy, defined as the fraction of subtasks performed by multiple agents ranges from 0.41 (centralized) to 0.50 (decentralized) for multi-agent systems (Table 5). The scaling law identifies a weak positive interaction with agent count ($\hat{\beta}_{R\times n_{a}}=0.041$, 95% CI: $[0.002,0.081]$, $p=0.040$), suggesting redundancy offers modest error-correction benefits when more agents participate. For a 4-agent system with $R=0.50$:
$\Delta P_{\text{redundancy}}=0.041\times 0.50\times 4=0.082,$
@@
$P_{\text{SA}}^{*}=\frac{\hat{\beta}_{4}}{\hat{\beta}_{17}}\approx\frac{0.063}{0.408}=0.154\quad\text{(in standardized units)},$
@@
+ Coordination metrics (Table 5) 0.589 0.513 190.3-190.3 20
@@
Table 4: Scaling principle coefficients relating performance to intelligence, task properties, and empirical coordination metrics ($R^{2}_{\text{train}}=0.589$, $R^{2}_{\text{CV}}=0.513$, $n=180$, AIC=190.3-190.3). Model uses 5-fold cross-validation; all significant predictors shown.
@@
Intelligence ($I$) -0.180 [-0.312, -0.048] 0.008 Linear capability effect
Intelligence2 ($I^{2}$) 0.256 [0.064, 0.449] 0.010 Accelerating returns to capability
$\log(1+T)$ 0.535 [0.347, 0.723] <<0.001 Tool diversity benefit
Single-Agent Baseline ($P_{\text{SA}}$) 0.319 [0.186, 0.453] <<0.001 Task difficulty proxy
@@
$P_{\text{SA}}\times\log(1+n_{a})$ 0.408-0.408 [-0.564, -0.251] <<0.001 Baseline paradox
$E_{c}\times T$ 0.330-0.330 [-0.432, -0.228] <<0.001 Efficiency-tools trade-off
$O\%\times T$ 0.141-0.141 [-0.213, -0.069] <<0.001 Overhead scales with task complexity
$A_{e}\times T$ 0.097-0.097 [-0.167, -0.027] 0.007 Error propagation in tool-rich systems
Coordination Structure
$\log(1+O\%)$ 0.032 [-0.004, 0.067] 0.084 Direct overhead cost (marginal)
$R\times n_{a}$ 0.041 [0.002, 0.081] 0.040 Redundancy benefit with scale
Table 5: Coordination metrics across architectures and families ($n=180$ configurations, 14,742 total instance runs). All systems matched for total reasoning tokens (mean $\mu=4,800$ per trial).
@@ 4.4 Coordination Efficiency, Error Dynamics, and Information Transfer
where $c$ is messages per reasoning turn. Performance plateaus near $c^{*}=0.39$ messages/turn (achieved by Decentralized and Centralized architectures at 0.41 and 0.39 respectively), corresponding to the success rates of 47.7% and 46.3%. Beyond this point, additional messages yield diminishing returns: Hybrid systems (515% coordination overhead, $T=44.3$) achieve only $\approx 2\%$–3% success gain versus Centralized (285% overhead, $T=27.7$), a difference of 0.7 percentage points that is not statistically significant ($t(178)=0.61$, $p=0.542$). This saturation reflects fundamental information limits in open-ended reasoning rather than mechanism failures: high-performing runs show convergent token overlap (shared tokens: mean $\approx 1.8$ bits; $p<0.001$ vs. low performers) suggesting message consensus is reached; further messages add redundancy rather than novel information.
@@
Figure 4: Agent Heterogeneity Effects on Multi-Agent Performance. Performance comparison of centralized (Orchestrator-Subagents) and decentralized (Peer Debate with Voting) multi-agent architectures on BrowseComp-Plus benchmark across three LLM families. High-capability models include GPT-5, Claude Sonnet 4.5, and Gemini-2.5 Pro; low-capability models include GPT-5 nano, Claude Sonnet 3.7, and Gemini-2.0 Flash. (1) Anthropic models uniquely benefit from heterogeneous mixing in centralized architecture, where low-capability orchestrator with high-capability subagents (0.42) outperforms homogeneous high-capability (0.32) by 31%, while OpenAI and Gemini show performance degradation under heterogeneous centralized configurations. (2) Decentralized mixed-capability approaches achieve near-optimal or superior performance compared to homogeneous high-capability baselines (OpenAI: 0.53 vs 0.50; Anthropic: 0.47 vs 0.37; Gemini: 0.42 vs 0.43), suggesting effective emergent collaboration despite capability asymmetry. (3) Centralized architectures with low-capability orchestrators underperform dramatically for OpenAI and Gemini families, indicating architectural constraints when coordination relies on less capable models.
@@
Architectural rankings remained stable across domains (Kendall $\tau=0.89$, coefficient of variation <0.1<0.1 across architectures), indicating coordination principles transcend specific task structures. Leave-one-domain-out cross-validation achieved $R^{2}=0.89$ ($p<0.001$), confirming that coordination effects generalize beyond the four benchmarks. Extrapolation to larger teams (n=6n=6–10) via the fitted power law yields 95% prediction intervals $3.2,6.8]\times$ turn-count increases (bootstrap coverage 94.2%), with high confidence in scaling behavior. Specifically, at $n=6$ agents, predicted turns range from 12.8 to 20.1 (SAS: 7.2; Centralized would reach $\approx 85$–130 turns). This super-linear scaling confirms the hard resource ceiling: beyond 3–4 agents, per-agent reasoning quality degrades sharply under fixed budgets.
@@
Cross-family analysis reveals distinct architectural preferences. OpenAI models show strongest Hybrid synergy on structured tasks (Finance: 52% success Hybrid vs. 39% SAS; Workbench: 56% Hybrid vs. 42% SAS). Anthropic models display most conservative, stable Centralized performance (mean 43% across tasks, SD = 2.3%, lowest variance). Google models exhibit robust cross-architecture efficiency (performance range < 5% across topologies). These patterns ($R^{2}=0.89$ cross-validation) reflect fundamental differences in attention mechanisms, activation sparsity, and representation geometry enabling or constraining multi-agent interaction, not superficial hyperparameter differences.

### v2

@@ 4.3 Scaling principles
Mixed-Effects Model Achieves 52.4% Cross-Validated Variance Explanation.
We fit a scaling principle to all 180 configurations that relates agentic system performance to four categories of predictors: 1) base model capability (intelligence index $I$), 2) system configuration (agent count $n_{a}$), 3) task properties (tool count $T$, single-agent baseline $P_{\text{SA}}$). These are instance-level predictors capturing within-benchmark variation, distinct from the benchmark-level domain complexity $D$ defined in Appendix C, and 4) empirically measured coordination metrics from Table 5 (efficiency $E_{c}$, overhead $O\%$, error amplification $A_{e}$, message density $c$, redundancy $R$). Rather than including all possible terms, we construct the model based on specific mechanistic hypotheses.
@@
$P=\beta_{0}$ $+{\color[rgb]{0,0.3984375,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3984375,0.80078125}\beta_{1}(I-\bar{I})}+\beta_{2}(I-\bar{I})^{2}+\beta_{3}\log(1+T)+\beta_{4}\log(1+n_{a})$
$+{\color[rgb]{0,0.3984375,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3984375,0.80078125}\beta_{5}\log(1+O\%)}+\beta_{6}c+\beta_{7}R+{\color[rgb]{0,0.3984375,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3984375,0.80078125}\beta_{8}E_{c}}+\beta_{9}\log(1+A_{e})$
$+{\color[rgb]{0,0.3984375,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.3984375,0.80078125}\beta_{10}P_{\text{SA}}}+\beta_{11}(I\times E_{c})+\beta_{12}(A_{e}\times P_{\text{SA}})$
$+{\color[rgb]{0.80078125,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0,0}\beta_{13}(O\%\times T)}+\beta_{14}(R\times n_{a})+\beta_{15}(c\times I)$
$+{\color[rgb]{0.80078125,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0,0}\beta_{16}(E_{c}\times T)}+{\color[rgb]{0.80078125,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0,0}\beta_{17}(P_{\text{SA}}\times\log(1+n_{a}))}$
@@
$\blacksquare$ Critical interactions ($p<0.001$)    $\blacksquare$ Significant effects ($p<0.05$)    $\blacksquare$ Non-significant ($p>0.05$)
where all predictors are standardized ($\mu=0$, $\sigma=1$) for interpretability. Log transformations are applied to right-skewed variables spanning multiple orders of magnitude ($O\%$: 0–515%; $T$: 4–16; $n_{a}$: 1–4; $A_{e}$: 1.0–17.2) to satisfy linearity assumptions. The $A_{e}\times T$ interaction retains $A_{e}$ without additional log transformation because $\log(1+A_{e})$ already appears as a main effect; including $\log(1+A_{e})\times T$ would introduce near-collinearity (VIF >8>8). Sensitivity analysis confirms qualitatively identical results under alternative specifications ($\Delta R^{2}_{\text{CV}}<0.01$). We validate model complexity through five-fold cross-validation with experiment-level holdout, which yields $R^{2}_{\text{CV}}=0.524$ ($\pm 0.033$ SD), mean absolute error MAE =0.089=0.089 ($\pm 0.011$), and root mean squared error RMSE =0.112=0.112 ($\pm 0.014$). The modest gap between training and cross-validated $R^{2}$ ($\Delta R^{2}=0.076$), combined with stable coefficient estimates across folds (coefficient of variation $18\%$ for all $|\hat{\beta}|>0.05$), indicates that the 20 parameters are justified by predictive power rather than overfitting. This model substantially outperforms simpler alternatives using only architectural labels ($R^{2}_{\text{CV}}=0.43$) or intelligence alone ($R^{2}_{\text{CV}}=0.28$), as shown in Table 3. Critically, this equation contains no dataset-specific parameters, enabling prediction on unseen task domains. Bootstrap resampling (n=1,000n=1{,}000 iterations) confirms coefficient stability (mean bootstrap SE <0.015<0.015 for all $|\hat{\beta}|>0.1$), and residual diagnostics satisfy normality (Shapiro–Wilk $p=0.412$) and homoscedasticity (Breusch–Pagan $p=0.298$), with residual standard error $\hat{\sigma}=0.118$. We evaluated regularized alternatives: Lasso (10-fold CV for $\lambda$ selection) retained 16 of 20 predictors with $R^{2}_{\text{CV}}=0.506$; Ridge achieved $R^{2}_{\text{CV}}=0.509$. Given minimal improvement and the interpretability benefits of the full model, we retain the unregularized specification.
The Efficiency-Tools Interaction Dominates Multi-Agent Performance ($\hat{\beta}=-0.267$, $p<0.001$).
@@
$\Delta P_{\text{efficiency}}=-0.267\times E_{c}\times T=\begin{cases}-1.99&\text{(single-agent, $E_{c}=0.466$)}\\ -0.32&\text{(multi-agent, $E_{c}=0.074$)}\end{cases}$
[…]
Error Amplification Exhibits Architecture-Dependent Catastrophic Failure Modes. Table 5 reveals dramatic variance in error amplification factors: single-agent ($A_{e}=1.0$), centralized ($A_{e}=4.4$), decentralized ($A_{e}=7.8$), hybrid ($A_{e}=5.1$), and strikingly, independent multi-agent ($A_{e}=17.2$). After controlling for other coordination metrics, neither the main effect of error amplification ($\hat{\beta}=-0.022$, $p=0.441$) nor its interaction with tool count ($A_{e}\times T$: $\hat{\beta}=-0.019$, $p=0.506$) reaches statistical significance. This suggests that the dramatic performance differences across architectures observed in Table 5 are better explained by other coordination mechanisms—particularly efficiency ($E_{c}$) and overhead ($O\%$)—rather than error propagation per se. Independent architecture’s universal underperformance (mean success 0.370 vs. 0.466 SAS) stems from absence of inter-agent communication: each agent operates in isolation, duplicating errors without correction opportunities, but this effect is subsumed by the efficiency metric ($E_{c}=0.234$ for Independent vs. $E_{c}=0.466$ for SAS).
@@
$O\%_{\text{max}}(T)=\frac{\hat{\beta}_{5}}{\hat{\beta}_{13}T}\log(1+O\%)\approx\frac{0.034}{0.162T}\log(1+O\%),$
beyond which overhead cost exceeds any coordination benefit. For $T=16$, this threshold is $O\%\approx 150\%$, ruling out all multi-agent architectures except possibly decentralized (263%, but compensated by parallelization). Empirically, workbench confirms this prediction: decentralized (mean 0.664) outperforms centralized (0.621) despite higher overhead, due to its superior parallel efficiency. This overhead-complexity interaction constitutes the third-strongest effect ($|\beta|=0.162$), reinforcing that coordination costs are not fixed but scale super-linearly with environmental complexity.
Intelligence Shows Linear Positive Effect ($\hat{\beta}_{I}=0.171$, $p=0.001$).
After centering intelligence scores to address multicollinearity (VIF reduced from 200 to 1.1), the linear capability effect becomes significant: higher-capability models achieve proportionally better performance across all architectures. The quadratic term ($I^{2}$) is not significant ($p=0.509$), indicating that capability scaling follows a linear rather than accelerating pattern within the tested range ($I\in[42,71]$). This finding suggests that coordination benefits scale consistently with model capability, without evidence of emergent super-linear gains at higher intelligence levels.
Redundancy Provides Marginal Benefit at Scale ($\hat{\beta}_{R\times n_{a}}=0.047$, $p=0.001$).
Work redundancy, defined as the fraction of subtasks performed by multiple agents ranges from 0.41 (centralized) to 0.50 (decentralized) for multi-agent systems (Table 5). The scaling law identifies a weak positive interaction with agent count ($\hat{\beta}_{R\times n_{a}}=0.047$, 95% CI: $[0.019,0.075]$, $p=0.001$), suggesting redundancy offers error-correction benefits when more agents participate. For a 4-agent system with $R=0.50$:
$\Delta P_{\text{redundancy}}=0.047\times 0.50\times 4=0.094,$
@@
$P_{\text{SA}}^{*}=-\frac{\hat{\beta}_{4}}{\hat{\beta}_{17}}\approx\frac{0.052}{0.404}=0.129\quad\text{(in standardized units)},$
@@
+ Coordination metrics (Table 5) 0.613 0.524 201.2-201.2 20
@@
Table 4: Complete scaling principle coefficients relating performance to intelligence, task properties, and empirical coordination metrics ($R^{2}_{\text{train}}=0.613$, $R^{2}_{\text{CV}}=0.524$, $n=180$, AIC=201.2-201.2). Intelligence is mean-centered ($\bar{I}=56.9$) to address multicollinearity between II and $I^{2}$ (VIF reduced from 200 to 1.1). Model uses 5-fold cross-validation. Non-significant terms ($p>0.05$) indicated with †.
@@
Intercept ($\beta_{0}$) 0.453 [0.433, 0.472] <<0.001 Baseline performance
Intelligence ($I-\bar{I}$) 0.171 [0.070, 0.272] 0.001 Linear capability effect
Intelligence2 ($(I-\bar{I})^{2}$) 0.007 [-0.013, 0.026] 0.509† Quadratic capability (not significant)
$\log(1+T)$ 0.411 [0.291, 0.531] <<0.001 Tool diversity benefit
$\log(1+n_{a})$ 0.052 [-0.061, 0.166] 0.367† Agent count effect
Single-Agent Baseline ($P_{\text{SA}}$) 0.315 [0.185, 0.445] <<0.001 Task difficulty proxy
Coordination Structure
$\log(1+O\%)$ 0.034 [0.011, 0.057] 0.005 Direct overhead cost
Message density ($c$) -0.057 [-0.110, -0.003] 0.039 Communication intensity
Redundancy ($R$) -0.007 [-0.052, 0.037] 0.748† Work overlap
Efficiency ($E_{c}$) -0.043 [-0.078, -0.007] 0.021 Coordination efficiency
$\log(1+A_{e})$ -0.022 [-0.077, 0.034] 0.441† Error amplification
@@
$P_{\text{SA}}\times\log(1+n_{a})$ -0.404 [-0.557, -0.252] <<0.001 Baseline paradox
$E_{c}\times T$ -0.267 [-0.355, -0.178] <<0.001 Efficiency-tools trade-off
$O\%\times T$ -0.162 [-0.241, -0.083] <<0.001 Overhead scales with task complexity
$A_{e}\times T$ -0.019 [-0.075, 0.037] 0.506† Error propagation in tool-rich systems
$R\times n_{a}$ 0.047 [0.019, 0.075] 0.001 Redundancy benefit with scale
$I\times E_{c}$ -0.022 [-0.075, 0.030] 0.404† Capability-efficiency
$A_{e}\times P_{\text{SA}}$ -0.065 [-0.146, 0.015] 0.114† Error-baseline
$c\times I$ -0.011 [-0.057, 0.034] 0.626† Communication-capability
$I\times\log(1+T)$ -0.069 [-0.138, 0.000] 0.053† Capability-tools
Table 5: Coordination metrics across architectures and families ($n=180$ configurations, 15,750 total instance runs). All systems matched for total reasoning tokens (mean $\mu=4,800$ per trial).
@@ 4.4 Coordination Efficiency, Error Dynamics, and Information Transfer
where $c$ is messages per reasoning turn. Performance plateaus near $c^{*}=0.39$ messages/turn (achieved by Decentralized and Centralized architectures at 0.41 and 0.39 respectively), corresponding to the success rates of 47.7% and 46.3%. Beyond this point, additional messages yield diminishing returns: Hybrid systems (515% coordination overhead, $T=44.3$) shows -2.4% versus Centralized (285% overhead, $T=27.7$), a difference of 1.1% that is not statistically significant ($t(178)=0.61$, $p=0.542$). This saturation reflects fundamental information limits in open-ended reasoning rather than mechanism failures: high-performing runs show convergent token overlap (shared tokens: mean $\approx 1.8$ bits; $p<0.001$ vs. low performers) suggesting message consensus is reached; further messages add redundancy rather than novel information.
@@
Figure 4: Agent Heterogeneity Effects on Multi-Agent Performance. Performance comparison of centralized (Orchestrator-Subagents) and decentralized (Peer Debate with Voting) multi-agent architectures on BrowseComp-Plus benchmark across three LLM families. High-capability models include GPT-5, Claude Sonnet 4.5, and Gemini-2.5 Pro; low-capability models include GPT-5 nano, Claude Sonnet 3.7, and Gemini-2.0 Flash. (1) Anthropic models uniquely benefit from heterogeneous mixing in centralized architecture, where low-capability orchestrator with high-capability subagents (0.42) outperforms homogeneous high-capability (0.32) by 31%, while OpenAI and Gemini show performance degradation under heterogeneous centralized configurations. (2) Decentralized mixed-capability approaches achieve near-optimal or superior performance compared to homogeneous high-capability baselines (OpenAI: 0.53 vs 0.50; Anthropic: 0.47 vs 0.37; Gemini: 0.42 vs 0.43), suggesting effective emergent collaboration despite capability asymmetry. (3) In centralized architectures, configurations with high-capability sub-agents outperform those with high-capability orchestrators across all model families, suggesting sub-agent capability matters more than orchestrator capability.
@@
Architectural rankings remained stable across domains (Kendall $\tau=0.89$, coefficient of variation <0.1<0.1 across architectures), indicating coordination principles transcend specific task structures. Extrapolation to larger teams (n=6n=6–10) via the fitted power law yields 95% prediction intervals $3.2,6.8]\times$ turn-count increases (bootstrap coverage 94.2%), with high confidence in scaling behavior. Specifically, at $n=6$ agents, predicted turns range from 12.8 to 20.1 (SAS: 7.2; Centralized would reach $\approx 85$–130 turns). This super-linear scaling confirms the hard resource ceiling: beyond 3–4 agents, per-agent reasoning quality degrades sharply under fixed budgets.
@@
Cross-family analysis reveals distinct architectural preferences. OpenAI models show strongest Hybrid synergy on structured tasks (Finance: 52% success Hybrid vs. 39% SAS; Workbench: 56% Hybrid vs. 42% SAS). Anthropic models display most conservative, stable Centralized performance (mean 43% across tasks, SD = 2.3%, lowest variance). Google models exhibit robust cross-architecture efficiency (performance range < 5% across topologies). These patterns reflect fundamental differences in attention mechanisms, activation sparsity, and representation geometry enabling or constraining multi-agent interaction, not superficial hyperparameter differences.